# CVOCSemRPL: Class-Variance Optimized Clustering, Semantic Information Injection and Restricted Pseudo Labeling based Improved Semi-Supervised Few-Shot Learning

### v2

@@ Abstract
Few-shot learning has been extensively explored to address problems where the amount of labeled samples is very limited for some classes. In the semi-supervised few-shot learning setting, substantial quantities of unlabeled samples are available. Such unlabeled samples are generally cheaper to obtain and can be used to improve the few-shot learning performance of the model. Some of the recent methods for this setting rely on clustering to generate pseudo-labels for the unlabeled samples. Since the effectiveness of clustering heavily influences the labeling of the unlabeled samples, it can significantly affect the few-shot learning performance. In this paper, we focus on improving the representation learned by the model in order to improve the clustering and, consequently, the model performance. We propose an approach for semi-supervised few-shot learning that performs a class-variance optimized clustering coupled with a cluster separation tuner in order to improve the effectiveness of clustering the labeled and unlabeled samples in this setting. It also optimizes the clustering-based pseudo-labeling process using a restricted pseudo-labeling approach and performs semantic information injection in order to improve the semi-supervised few-shot learning performance of the model. We experimentally demonstrate that our proposed approach significantly outperforms recent state-of-the-art methods on the benchmark datasets.
@@ 1 Introduction
Modern deep networks often surpass human performance but demand ever deeper architectures and vast labeled datasets, whose annotation is costly and time-consuming. This data bottleneck motivates few-shot learning, where models must generalize from only a handful of examples. Therefore, researchers have been looking at few-shot learning (FSL) as a solution. Few-shot learning tackles new classes with only a handful of labeled examples by first training a feature extractor on richly annotated base classes and then transferring those learned representations to extract meaningful embeddings for the scarce few-shot samples. The performance of existing FSL methods is still limited due to the extremely limited availability of labeled samples [1, 2]. Unlabeled data is often plentiful and easy to collect, so incorporating large pools of unlabeled examples alongside the scarce few-shot samples, known as semi-supervised few-shot learning, can substantially boost performance. In this paper, we tackle the semi-supervised few-shot learning (SSFL) problem.
Researchers have proposed multiple approaches to address the SSFSL problem, such as clustering to assign pseudo-labels to unlabeled samples and utilizing them to improve the FSL performance [3, 4]. Clustering relies heavily on the quality of representation learned by the network, especially on the base classes, since this ensures that newly observed images, even from unseen classes, are far away from images of other classes and close to images from the same class in the representation space. A model with better quality representation will facilitate better clustering of extracted features and consequently lead to better pseudo-label assignment to the unlabeled samples (refer to Table 5).
We propose a novel SSFSL approach in this work. Specifically, we use a class variance optimized clustering (CVOC) process to cluster given images and use a semantic injection network to refine the cluster centers by incorporating semantic information. The proposed CVOC process optimizes intra- and inter-class separation and utilizes a cluster separation tuner to achieve improved clustering (refer to Table 5). We use the refined cluster centers to perform restricted pseudo-labeling of the unlabeled samples and incorporate all labeled samples to make predictions for the query samples (see Sec. 4). We perform experiments on multiple benchmark datasets and demonstrate that our proposed approach significantly outperforms the compared approaches.
@@
    We experimentally demonstrate how incorporating semantic features of classes and managing the intra-class compactness and inter-class separation helps to further improve the performance of the model, as shown in Table 5.
@@ 2.1 Few-shot Image Classification
Few-shot learning (FSL) methodologies [5, 6, 7, 8, 9] generally pre-train the model using extensive data from the base classes and subsequently modify the pre-trained model for recognizing novel categories.  [10] utilizes an average feature prototype of the classes with very few samples and performs predictions by identifying the nearest prototype. TADAM [11] adapts the features of the few-shot class images using the few support images available for such classes. [12] improves upon prototypical network using an importance-weighting strategy for the different support samples of a class and makes the prototypes more informative. Meta-learning methods learn to quickly adapt to new classes using very few samples, and these methods involve training a model at a specific level as well as at an abstract level, referred to as learning to learn.  [13]. Researchers have proposed various meta-learning-based FSL approaches, such as  [14, 15].
@@ 2.2 Pseudo-labeling based Semi-supervised Few-shot Methods
The semi-supervised few-shot learning (FSL) setting is a modified FSL setting with access to unlabeled samples. Several semi-supervised FSL methods assign pseudo-labels to unlabeled samples and then consider these samples as labeled samples for the few-shot classification task [16, 17, 13]. However, if some of the pseudo labels are incorrect, the performance of the model is degraded. Some methods utilize label propagation to generate pseudo-labels [18][19], selectively sample high-confidence labels, but treat each sample independently and miss out on broader relational context.  [20] attempts to unify pseudo-labeled data in a single metric space, but also does not consider the inter-sample relationships. Poisson Transfer Network (PTN) [21] incorporates some relational understanding between labeled and unlabeled samples. However, it lacks the capacity to effectively mine inter- and intra-class relationships. Instance Credibility Inference (ICI) [22] ranks pseudo-labels based on credibility and sparsity, but neglects the issue of low-quality pseudo-labels. Cluster-FSL [4] employs a multi-factor clustering (MFC) strategy, which synthesizes various information sources from labeled data to produce both soft and hard pseudo-labels. It pseudo-labels selected unlabeled samples. However, we experimentally observed that the pseudo-labeling strategy employed by it led to most of the unlabeled images becoming eligible for pseudo-labeling, which would degrade the clustering process with outlier samples or incorrectly pseudo-labeled samples. Cluster-FSL’s limitations are rooted in its inadequate handling of inter-class and intra-class relationships, which restricts its ability to differentiate effectively between distinct classes and to capture subtle variations within individual classes. This leads to unlabeled samples with incorrect pseudo-labels being considered as labeled samples and consequently limiting the model performance.
In this work, we propose to mitigate these limitations through a class-variance optimized clustering (CVOC) approach and, which aligns the embedding space according to both inter- and intra-class relationships using targeted loss functions and also by using cluster separation tuner (CST). The proposed approach also incorporates semantic features to further improve the clustering process. Further, the proposed approach performs restricted pseudo-labeling and allocates pseudo-labels exclusively to unlabeled samples with low entropy predictions, reducing the possibility of noisy pseudo-labels affecting the model performance.
@@ 3 Problem Setting
In the few-shot learning (FSL) setting, there are two sets of non-overlapping classes: the base classes $C_{b}$ and the novel/few-shot classes $C_{f}$. The base classes have a sufficient number of training samples, while the novel classes have very few training samples. The dataset is split into the train set $D_{tr}$, the validation set $D_{va}$, and the test set $D_{tt}$. The train set $D_{tr}$ only contains samples from the base classes, while the validation $D_{va}$ and test $D_{tt}$ sets contain samples from the novel classes without any overlap between the classes in $D_{va}$ and $D_{tt}$. Data is provided as $N$-way $K$-shot episodes, where $N$ refers to the number of classes in the episode, and $K$ refers to the number of support/training samples from each of these classes. The support set can be denoted as $S={(x^{s}_{i},y^{s}_{i})}^{N\times K}_{i=0}$. The episode also contains $T$ randomly selected query/test samples from each of these classes on which the model can be evaluated. The query set can be denoted as $Q={(x^{q}_{i},y^{q}_{i})}^{N\times T}_{i=0}$. The episodes from the train set contain a support set $S$ and a query set $Q$ from $N$ randomly selected classes from among the classes in the train set. Similarly, the episodes from the test set contain a support set $S$ and a query set $Q$ from $N$ randomly selected classes in the test set. In this paper, we address the semi-supervised FSL setting. In this setting, apart from the above episodic data, we also have access to $u$ randomly selected samples from the classes in the episode that are not labeled. The unlabeled set can be denoted as $U={(x_{i}^{a})}^{N\times u}_{i=0}$. This leads to a setting where we have very few training samples for the classes in the episode but a good quantity of unlabeled samples.
@@ 4.1 Components
The proposed approach involves training a visual feature extractor network ($M$) and a semantic injection network ($S$). The visual feature extractor $M$ is a simple deep neural network that, once trained, should be able to extract good features from images. The semantic injection network contains an encoder network ($E_{s}$) and a decoder network ($D_{s}$). Once trained, the semantic injection network $S$ should be able to map features extracted by the trained visual feature extractor from an image concatenated with the semantic feature for the corresponding class to a joint embedding space where the joint feature embedding is close to the image feature prototype of the same class.
@@ 4.2 Pre-training the Visual Feature Extractor
The visual feature extractor $M$ is first pre-trained on the train set $D_{tr}$, containing the base classes, in a regular batch-wise supervised setting. Following existing methods, we train $M$ simultaneously on the image label classification task and the rotation prediction task (see Eq. 1), since this joint training has been shown to improve the generalization power of the network. We add an image label classification head $F_{c}$ and a rotation angle classification head $F_{r}$ to $M$ for this training process.
@@
Where, $y_{i}$ and $r_{i}$ refer to the image label and rotation label for image $x_{i}$, respectively. $L_{ce}$ refers to the cross-entropy loss.
@@ 4.3 Finetuning the Visual Feature Extractor with Class-Variance Optimized Clustering
The pre-trained visual feature extractor $M$ is fine-tuned on episodes from the train set $D_{tr}$. In the few-shot regime, with only a handful of labeled support examples in each episode, aggressive augmentations greatly increase input diversity and prevent the model from simply memorizing the small support set. Following recent SSFSL methods [4], we begin by applying RandAugment [23] to every support image to guard against over-fitting. After extracting embeddings for both these augmented support samples and the query set with our visual feature extractor $M$, we refine all features using embedding propagation [3, 4], a technique commonly used in SSFSL to obtain manifold-smoothed embeddings. Finally, the propagated representations are fed into our Class-Variance Optimized Clustering (CVOC) module (Fig. 1). Then, based on the support set, the CVOC module and label propagation (LP), obtain the soft labels of the query set samples. Label propagation [24] is a graph-based technique commonly used in SSFSL to “propagate" labels from labeled samples to unlabeled samples, utilizing the similarity between the samples as weights for edges between the graph nodes.
@@ 4.3.1 Class-Variance Optimized Clustering and Cluster Separation Tuner
The proposed class-variance optimized clustering procedure utilizes an improved clustering procedure and a cluster separation tuner (CST) to get a more robust clustering of the labeled and unlabeled image features in the episode. For every class label $c\in\{1,\dots,C\}$ we build a factor dictionary
$\mathbf{F}_{c}=\Bigl{[}\underbrace{\mathbf{s}^{(c)}_{1},\dots,\mathbf{s}^{(c)}_{K}}_{\text{support}},\;\underbrace{P_{0}(c)}_{\text{initial prototype}}\Bigr{]}\in\mathbb{R}^{d\times(K+1)},$ (2)
where $\mathbf{s}^{(c)}_{k}\in\mathbb{R}^{d}$ is the $k$-th support embedding of class $c$ and $P_{0}(c)$ is their mean. Given a feature vector $x$, its representation weights (solution in the collaborative representation [25]) and reconstruction distance are
$\beta^{\!*}_{c}(x)=\arg\min_{\beta}\bigl{\|}x-\mathbf{F}_{c}\beta\bigr{\|}_{2}^{2}+\lambda\|\beta\|_{2}^{2},\quad d_{\text{rec}}(x,c)=\bigl{\|}x-\mathbf{F}_{c}\beta^{\!*}_{c}(x)\bigr{\|}_{2}^{2},$ (3)
The reconstruction-based distance captures the true shape and spread of each category by leveraging both the cluster center and representative labeled samples. This richer, multi-prototype view helps us account for within-category heterogeneity and yields more accurate, robust assignments for unlabeled data; compared to traditional clustering techniques. But $d_{\text{rec}}$ alone may be ambiguous when class subspaces overlap, so we regularise it with two Euclidean terms. Let $P(c)$ denote the current prototype of class $c$ (updated after every assignment step) and $\operatorname{ep}\!\bigl{(}M(x)\bigr{)}$ the embedding–propagated feature of $x$ as in [3]. Using $\mathcal{S}$ to denote the support set (expanded support set through augmentation), we define
@@
Here $C$ is the number of classes in the episode. The intra-class term $L_{\text{intra}}$ computes, for every support sample, the squared Euclidean distance between its embedding (after embedding propagation) and the prototype of its own class, and then averages these distances over the entire support set. Conversely, the inter-class term $L_{\text{inter}}$ takes each support sample, measures its squared distance to every other class prototype, averages those distances across the $C{-}1$ “foreign” classes, and finally averages the resulting values over all support samples.
The distance used both for cluster assignment and for the logits that supervise $M$ is computed as follows.
@@
with weights $w_{\text{intra}},w_{\text{inter}}$ balancing the pull–push forces. Basically, the regularizer terms shrink intra-class spread and enlarge prototype margins, leading to cleaner pseudo labels and superior few-shot accuracy relative to the previous methods (see Table 5).
Each sample $x$ (both support and query) is (re)labeled with the class that minimizes the above distance. After that, every prototype is reset to the mean of its current members. Next, we refine this set with a lightweight cluster separation tuner (CST) modelled after the continuous firefly heuristic [26]. At every CST iteration, we compute a brightness score for each prototype.
@@
where the first two terms reuse the same intra- and inter-class distances, but taken classwise rather than episode-wise here, and the last term penalises prototypes that sit closer than an $\varepsilon$ margin to any neighbour. If prototype $P(c)$ is dimmer than $P(c^{\prime})$ ($B(c)<B(c^{\prime})$) we move it towards the brighter neighbour with an attractiveness
@@
Here $\beta_{0}$ sets the maximum attraction, $\gamma$ controls its spatial decay, and the Gaussian–type exploration term of amplitude $\alpha$ (decayed each loop by $0.995$) prevents premature convergence. Iterating this procedure for a small, fixed number of loops maximizes prototype separation while keeping each class center tethered to its own support cloud. For additional details, implementation specifics, and related experiments, please refer to the supplementary materials.
To predict labels for each query embedding $x_{q}$, we first compute its per-class reconstruction residuals $d_{\mathrm{rec}}(x_{q},c)$ as in Eq. (3). These distances are then converted into logits as follows.
@@
Finally the predicted label is obtained as follows
$\hat{y}_{q}\;=\;\arg\max_{c}\,p(c\mid x_{q})\,.$ (12)

### v3

# Semantic-Anchored, Class Variance-Optimized Clustering for Robust Semi-Supervised Few-Shot Learning

@@ Abstract
Few-shot learning has been extensively explored to address problems where the amount of labeled samples is very limited for some classes. In the semi-supervised few-shot learning setting, substantial quantities of unlabeled samples are available. Such unlabeled samples are generally cheaper to obtain and can be used to improve the few-shot learning performance of the model. Some of the recent methods for this setting rely on clustering to generate pseudo-labels for the unlabeled samples. Since the effectiveness of clustering heavily influences the labeling of the unlabeled samples, it can significantly affect the few-shot learning performance. In this paper, we focus on improving the representation learned by the model in order to improve the clustering and, consequently, the model performance. We propose an approach for semi-supervised few-shot learning that performs a class-variance optimized clustering coupled with a cluster separation tuner in order to improve the effectiveness of clustering the labeled and unlabeled samples in this setting. It also optimizes the clustering-based pseudo-labeling process using a restricted pseudo-labeling approach and performs semantic information injection in order to improve the semi-supervised few-shot learning performance of the model. We experimentally demonstrate that our proposed approach significantly outperforms recent state-of-the-art methods on the benchmark datasets. To further establish its robustness, we conduct extensive experiments under challenging conditions, showing that the model generalizes well to domain shifts and achieves new state-of-the-art performance in open-set settings with distractor classes, highlighting its effectiveness for real-world applications.
@@ 1 Introduction
Modern deep networks often surpass human performance but demand ever deeper architectures and vast labeled datasets, whose annotation is costly and time‐consuming. This data bottleneck motivates few‐shot learning, where models must generalize from only a handful of examples. Therefore, researchers have been looking at few-shot learning (FSL) as a solution. Few‐shot learning tackles new classes with only a handful of labeled examples by first training a feature extractor on richly annotated base classes and then transferring those learned representations to extract meaningful embeddings for the scarce few‐shot samples. The performance of existing FSL methods is still limited due to the extremely limited availability of labeled samples[dong2024pseudo, sun2019meta]. Unlabeled data is often plentiful and easy to collect, so incorporating large pools of unlabeled examples alongside the scarce few-shot samples, known as semi-supervised few-shot learning, can substantially boost performance. In this paper, we tackle the semi-supervised few-shot learning (SSFL) problem.
Researchers have proposed multiple approaches to address the SSFSL problem, such as clustering to assign pseudo-labels to unlabeled samples and utilizing them to improve the FSL performance [rodriguez2020embedding, ling2022semi]. Clustering relies heavily on the quality of representation learned by the network, especially on the base classes, since this ensures that newly observed images, even from unseen classes, are far away from images of other classes and close to images from the same class in the representation space. A model with better quality representation will facilitate better clustering of extracted features and consequently lead to better pseudo-label assignment to the unlabeled samples (refer to Table 5).
We propose a novel SSFSL approach in this work. Specifically, we use a class variance optimized clustering (CVOC) process to cluster given images and use a semantic injection network to refine the cluster centers during inference by incorporating semantic information derived from the class labels. The proposed CVOC process optimizes intra- and inter-class separation and utilizes a cluster separation tuner to achieve improved clustering (refer to Table 5). We use the refined cluster centers to perform restricted pseudo-labeling of the unlabeled samples and incorporate all labeled samples to make predictions for the query samples (see Sec. 4). We perform experiments on multiple benchmark datasets and demonstrate that our proposed approach significantly outperforms the compared approaches.
@@
    We experimentally demonstrate how incorporating semantic features of classes and managing the intra-class compactness and inter-class separation leads to better feature representations in the embedding space, thereby improving the performance of the model, as shown in Table 5. We also provide a novel class description generation pipeline (see Section 8.4.4), which overcomes raw class name noise, poor label description,etc.
  4. 4.
    Furthermore, we validate the robustness and generalization capabilities of our approach under challenging realistic scenarios, including new state-of-the-art performance in distractive semi-supervised few-shot learning settings and strong cross-domain few-shot learning results. Refer to Section 8.1.
@@ 2.1 Few-shot Image Classification
Few-shot learning (FSL) methodologies [kozerawski2018clear, yoo2018efficient, rusu2018meta, gidaris2019generating, lee2019meta] generally pre-train the model using extensive data from the base classes and subsequently modify the pre-trained model for recognizing novel categories. [snell2017prototypical] utilizes an average feature prototype of the classes with very few samples and performs predictions by identifying the nearest prototype. TADAM [oreshkin2018tadam] adapts the features of the few-shot class images using the few support images available for such classes. [JI202113] improves upon prototypical network using an importance-weighting strategy for the different support samples of a class and makes the prototypes more informative. Meta-learning methods learn to quickly adapt to new classes using very few samples, and these methods involve training a model at a specific level as well as at an abstract level, referred to as learning to learn. [yu2020transmatch]. Researchers have proposed various meta-learning-based FSL approaches, such as [finn2017model, li2017meta].
@@ 2.2 Pseudo-labeling based Semi-supervised Few-shot Methods
The semi-supervised few-shot learning (FSL) setting is a modified FSL setting with access to unlabeled samples. Several semi-supervised FSL methods assign pseudo-labels to unlabeled samples and then consider these samples as labeled samples for the few-shot classification task [lazarou2021iterative, wu2018exploit, yu2020transmatch]. However, if some of the pseudo labels are incorrect, the performance of the model is degraded. Some methods utilize label propagation to generate pseudo-labels [liu2018learning]. [li2019learning] selectively sample high-confidence labels, but treat each sample independently and miss out on broader relational context. [huang2021pseudo] attempts to unify pseudo-labeled data in a single metric space, but also does not consider the inter-sample relationships. Poisson Transfer Network (PTN) [huang2021ptn] incorporates some relational understanding between labeled and unlabeled samples. However, it lacks the capacity to effectively mine inter- and intra-class relationships. Instance Credibility Inference (ICI) [wang2020instance] ranks pseudo-labels based on credibility and sparsity, but neglects the issue of low-quality pseudo-labels. Cluster-FSL [ling2022semi] employs a multi-factor clustering (MFC) strategy, which synthesizes various information sources from labeled data to produce both soft and hard pseudo-labels. It pseudo-labels selected unlabeled samples. However, we experimentally observed that the pseudo-labeling strategy employed by it led to most of the unlabeled images becoming eligible for pseudo-labeling, which would degrade the clustering process with outlier samples or incorrectly pseudo-labeled samples. Cluster-FSL’s limitations are rooted in its inadequate handling of inter-class and intra-class relationships, which restricts its ability to differentiate effectively between distinct classes and to capture subtle variations within individual classes. This leads to unlabeled samples with incorrect pseudo-labels being considered as labeled samples and consequently limiting the model performance.
In this work, we address these limitations through a dual-novelty framework. First, we introduce a class-variance optimized clustering (CVOC) approach, together with a cluster separation tuner (CST), which acts as a geometric regularizer that aligns the embedding space according to both inter- and intra-class relationships via targeted loss functions. This design yields a cleaner manifold that reduces overlap between classes and pulls in-class samples closer, which is critical for robustness under open-set distractors and domain shift. Second, we propose a Semantic Injection Module that learns a shared non-linear latent manifold between visual and textual features, going beyond standard linear projectors that suffer from hubness and misalignment. This module is model-agnostic and pre-trained separately on text and image features, so it can be plugged into any SS-FSL backbone without interfering with few-shot training, enabling test-time semantic alignment in a zero-leakage fashion. Our approach also incorporates high-quality semantic class descriptions and restricted pseudo-labeling, allocating pseudo-labels exclusively to unlabeled samples with low-entropy predictions, thereby reducing noisy labels while keeping inference latency low. Taken together, these components not only achieve state-of-the-art pseudo-label accuracy among clustering-based SS-FSL methods, but also position our framework as a forward-looking solution for realistic scenarios dominated by distractors and domain shifts, rather than idealized clean benchmarks.
@@ 3 Problem Setting
In the few-shot learning (FSL) setting, there are two sets of non-overlapping classes: the base classes $C_{b}$ and the novel/few-shot classes $C_{f}$. The base classes have a sufficient number of training samples, while the novel classes have very few training samples. The dataset is split into the train set $D_{tr}$, the validation set $D_{va}$, and the test set $D_{tt}$. The train set $D_{tr}$ only contains samples from the base classes, while the validation $D_{va}$ and test $D_{tt}$ sets contain samples from the novel classes without any overlap between the classes in $D_{va}$ and $D_{tt}$. Data is provided as $N$-way $K$-shot episodes/tasks, where $N$ refers to the number of classes in the episode, and $K$ refers to the number of support/training samples from each of these classes. The support set can be denoted as $S={(x^{s}_{i},y^{s}_{i})}^{N\times K}_{i=0}$. The episode also contains $T$ randomly selected query/test samples from each of these classes on which the model can be evaluated. The query set can be denoted as $Q={(x^{q}_{i},y^{q}_{i})}^{N\times T}_{i=0}$. The episodes from the train set contain a support set $S$ and a query set $Q$ from $N$ randomly selected classes from among the classes in the train set. Similarly, the episodes from the test set contain a support set $S$ and a query set $Q$ from $N$ randomly selected classes in the test set. In this paper, we address the semi-supervised FSL setting. In this setting, apart from the above episodic data, we also have access to $u$ randomly selected samples from the classes in the episode that are not labeled. The unlabeled set can be denoted as $U={(x_{i}^{a})}^{N\times u}_{i=0}$. This leads to a setting where we have very few training samples for the classes in the episode, but a good quantity of unlabeled samples.
@@ 4.1 Components
The proposed approach involves training a visual feature extractor network ($M$) and a semantic injection network ($S$). The visual feature extractor $M$ is a simple deep neural network that takes an image as input, and once trained, it should be able to extract good features from images. The semantic injection network contains an encoder module ($E_{s}$) and a decoder module ($D_{s}$). The encoder module $E_{s}$ takes as input the prototype feature of a class concatenated with the semantic feature of the same class and maps it to a joint embedding space, while the decoder module $D_{s}$ reconstructs the input to the encoder module from the joint embedding. Once trained, the semantic injection network $S$ should be able to map the input features to a joint embedding space where the joint feature embedding is close to the image feature prototype of the same class.
@@ 4.2 Pre-training the Visual Feature Extractor
The visual feature extractor $M$ is first pre-trained on the train set $D_{tr}$, containing the base classes, in a batch-wise supervised setting. Following various existing methods, we train $M$ simultaneously on the image label classification task and the rotation prediction task (see Eq. 1), since this joint training has been shown to improve the generalization power of the network. We add an image label classification head $F_{c}$ and a rotation angle classification head $F_{r}$ to $M$ for this training process. The images are rotated by four fixed angles and the model predicts the class label using $F_{c}$ and the rotation angle label using $F_{r}$. The pretraining loss $L_{pt}$ is defined below:
where, $y_{i}$ and $r_{i}$ refer to the image label and rotation label for image $x_{i}$, respectively. $L_{\mathrm{ce}}$ refers to the cross-entropy loss.
@@ 4.3 Finetuning the Visual Feature Extractor with Class-Variance Optimized Clustering
After pre-training the visual feature extractor $M$, we fine-tune it on episodes from the train set $D_{tr}$. In the few-shot regime, with only a handful of labeled support examples in each episode, aggressive augmentations greatly increase input diversity and prevent the model from simply memorizing the small support set. Following recent SSFSL methods [ling2022semi], we begin by applying RandAugment [cubuk2020randaugment] to every support image to guard against over-fitting. After extracting embeddings for both these augmented support samples and the query set with our visual feature extractor $M$, we refine all features using embedding propagation [rodriguez2020embedding, ling2022semi], a technique commonly used in SSFSL to obtain manifold-smoothed embeddings. This process is described in 8.4.1 in the Appendix. Finally, the propagated representations are fed into our Class-Variance Optimized Clustering (CVOC) module (Fig. 1). Then, based on the support set, the CVOC module, and label propagation (LP), we obtain the soft labels of the query set samples. Label propagation [iscen2019label] is a graph-based technique commonly used in SSFSL to “propagate" labels from labeled samples to unlabeled samples, utilizing the similarity between the samples as weights for edges between the graph nodes. It is also discussed in 8.4.1 in the Appendix.
@@ 4.3.1 Class-Variance Optimized Clustering and Cluster Separation Tuner
The proposed class-variance optimized clustering procedure utilizes an improved clustering procedure and a cluster separation tuner (CST) to get a more robust clustering of the labeled and unlabeled image features in the episode. For each class label $c\in\{1,\ldots,C\}$, we construct a class-specific factor dictionary $\mathbf{F}_{c}$ consisting of all support embeddings for that class together with its initial prototype, defined as the mean embedding of the corresponding support samples:
$\mathbf{F}_{c}=\big[s^{(c)}_{1},\ldots,s^{(c)}_{K},\;P_{0}(c)\big]\in\mathbb{R}^{d\times(K+1)},$
where $s^{(c)}_{k}\in\mathbb{R}^{d}$ denotes the $k$-th support embedding and $P_{0}(c)$ is the initial prototype. Given a feature vector $x$, we compute its coefficient vector by solving the regularized least-squares problem:
$\beta_{c}^{*}(x)=\arg\min_{\beta}\;\|x-\mathbf{F}_{c}\beta\|_{2}^{2}+\lambda\|\beta\|_{2}^{2},$
with $\lambda=0.01$. The reconstruction distance is then
$d_{\text{rec}}(x,c)=\|x-\mathbf{F}_{c}\beta_{c}^{*}(x)\|_{2}^{2},$
which measures how well $x$ can be represented using the support embeddings and prototype of class $c$.
Furthermore, the coefficient vector decomposes as
$\beta_{c}^{*}(x)=\big[\beta_{c,1}^{*}(x),\beta_{c,2}^{*}(x),\ldots,\beta_{c,N}^{*}(x)\big]^{\top},$
where each sub-vector $\beta_{c,j}^{*}(x)\in\mathbb{R}^{(K+1)\times 1}$ corresponds to the $j$-th cluster within the class, offering a richer representation of intra-class variability by allocating different coefficient subsets to different clusters.
The reconstruction-based distance captures the true shape and spread of each category by leveraging both the cluster center and representative labeled samples. This richer, multi-prototype view helps us account for within-category heterogeneity and yields more accurate, robust assignments for unlabeled data compared to traditional clustering techniques. However, $d_{\text{rec}}$ alone may be ambiguous when class subspaces overlap. Therefore, we regularize it with two Euclidean terms, i.e., $L_{\text{intra}}$ and $L_{\text{inter}}$. Let $P(c)$ denote the current prototype of class $c$ (updated after every assignment step) and $\operatorname{ep}\!\bigl(M(x)\bigr)$ denote the embedding propagated feature of $x$ as in [rodriguez2020embedding]. Using $\mathcal{S}$ to denote the support set (expanded support set obtained through augmentation), we define
@@
Here $C$ is the number of classes in the episode. Considering an image $x_{i}$ and label $y_{i}$, P($y_{i}$) denotes the prototype of the class $y_{i}$, and $P(j)$ denotes the prototype of some other class $j$ present in the episode. The intra-class term $L_{\text{intra}}$ computes, for every support sample, the squared Euclidean distance between its embedding (after embedding propagation) and the prototype of its own class, and then averages these distances over the entire support set. Conversely, the inter-class term $L_{\text{inter}}$ takes each support sample, measures its squared distance to every other class prototype, averages those distances across the C1C{-}1 “foreign” classes, and finally averages the resulting values over all support samples.
The combined distance term $d(x,c)$ utilizes $L_{\text{intra}}$ and $L_{\text{inter}}$ to regularize the reconstruction distance $d_{\text{rec}}$ as shown in Eq. 4. This distance is used both for cluster assignment and for the logits that supervise $M$.
@@
with weights $w_{\text{intra}},w_{\text{inter}}$ balancing the pull–push forces and $d_{rec}$ is the reconstruction-based distance between a feature vector $x$ and class $c$ representation. Basically, the regularizer terms shrink intra-class spread and enlarge prototype margins, leading to cleaner pseudo labels and superior few-shot accuracy relative to the previous methods (see Table 5).
In each episode, both support and query samples $x$ are (re)labeled with the class that minimizes the above distance. After that, every prototype is reset to the mean of its current members. Next, we refine the prototypes with a lightweight cluster separation tuner (CST) modelled after the continuous firefly heuristic [yang2013firefly]. At every CST iteration, we compute a brightness score $B(c)$ for the prototype of each class c.
@@
where the first two terms reuse the same intra- and inter-class distances, but taken class-wise(the intra-class and inter-class distances are computed separately for each individual class rather than averaged over all classes within the episode) rather than episode-wise here, and the last term penalizes prototypes that sit closer than an $\varepsilon$ margin to any neighbour. If prototype $P(c)$ is dimmer than $P(c^{\prime})$ ($B(c)<B(c^{\prime})$) we move it towards the brighter neighbour with an attractiveness
@@
Here $\beta_{0}$ sets the maximum attraction, $\gamma$ controls its spatial decay, and the Gaussian–type exploration term of amplitude $\alpha$ (decayed each loop by $0.995$) prevents premature convergence. Iterating this procedure for a small, fixed number of loops maximizes prototype separation while keeping each class center tethered to its own support cloud. Our method synthesizes the complementary strengths of reconstruction-based and Euclidean-based metrics. The reconstruction residual captures the intrinsic local geometry, while the Euclidean regularizer enforces global separability. The Cluster Separation Tuner (CST) further refines this by explicitly extending decision boundaries to maximize inter-class margins. For additional details, implementation specifics, and related experiments, please refer to Sec. 8.3 in the Appendix.
To predict labels for each query embedding $x_{q}$, we first compute its per‐class reconstruction residuals $d_{\mathrm{rec}}(x_{q},c)$. These distances are then converted into logits as follows.
@@
Only reconstruction residuals are used at inference because the variance optimization terms ($L_{\text{intra}},L_{\text{inter}}$) function as global structural constraints during the iterative transductive tuning phase to refine the class prototypes and separate clusters. However, the equation just above represents the final inference step after these prototypes have converged. At this stage, because the prototypes are already spatially optimized to account for class variance, we rely solely on the reconstruction residual to provide the most precise, instance-level measure of fit.
@@
Finally, the predicted label is obtained as: $\hat{y}_{q}\;=\;\arg\max_{c}\,p(c\mid x_{q})$